The area of the circle that can be inscribed in a square of side 6 cm in
Answer by addingup(3677)

You didn't complete your question, so I'll guess:
1) the area of the square is 6*6 = 36
2) the area of the circle is a = Pir^2 and the radius is 6/2 = 3, so the area is:
Pi*3^2 = Pi*9 = 28.27
36-28.27 = 7.73 is how much bigger the area of the square is compared to the area of the circle.
If this was not the question feel free to email me.
:
John